OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the right triangle with the given measures. c is the hypotenuse. a = 6, b = 10, c = ?

OpenStudy (anonymous):

\[c ^{2}=a ^{2}+b ^{2}\rightarrow c ^{2}=6^{2}+10^{2}=36+100=136\rightarrow c=+\sqrt{136}=2 sqrt{34}\]

OpenStudy (anonymous):

Use : \[c^2 = a^2 + b^2\] because In a right angled triangle, \[Hypotenuse^2 = Base^2 + Height^2\] Or you can use directly: \[c = \sqrt{a^2 + b^2}\] You have a = 6, you have b = 10 and now can you find c??